# Conceptualization of this study, Methodology, Software

### v1

@@ 1 Introduction
Environmental sound classification (ESC) has emerged as a crucial task in machine learning and signal processing, finding applications in diverse domains such as surveillance, smart city monitoring, fault detection in automobiles, and acoustic-based quality control in manufacturing. Unlike structured audio signals such as speech and music, environmental sounds exhibit significant variations in temporal and spectral characteristics, making classification a challenging problem. Recent advances in machine learning and deep learning have transformed the way environmental audio is analyzed Purwins2019. First, traditional machine learning approaches relied on handcrafted features such as Mel-Frequency Cepstral Coefficients (MFCCs), Zero-Crossing Rate (ZCR), and Spectral Contrast (SPC), among others, and their combinations chachada2014environmental; Bountourakis2015. They were used as inputs to classifiers like Support Vector Machines (SVMs), k-Nearest Neighbors (k-NN), Decision Trees (DT), Local Binary Pattern (LPB), and Random Forest (RF) Roma2018; Toffa2021. These models achieved reasonable results but often lack robustness to noise and generalization across diverse acoustic environments. With the advent of deep learning, particularly Convolutional Neural Networks (CNNs), ESC systems became capable of learning hierarchical time-frequency patterns directly from the spectrogram representations Purwins2019. They can efficiently extract local spectrum features, making them well-suited for multiple audio classification tasks hershey2017cnn; Zheng2018; Ghosal2018.
@@
However, CNNs often struggle to capture long-range dependencies in audio sequences, which is a fundamental limitation in modeling complex environmental sounds. In contrast, transformer-based models, such as AST  gong2021ast, leverage self-attention mechanisms to model global dependencies in spectrogram representations. ASTs have demonstrated state-of-the-art performance in various audio classification tasks, particularly when pre-trained on large-scale datasets gong2021ast. However, their reliance on extensive data and computational resources poses challenges for real-world applications where labeled data are limited. Within this broader evolution, researchers have also explored feature aggregation as a complementary approach to improve learning efficiency Liu2021. By combining multiple extracted features, researchers have sought to improve classification performance and reduce model complexity, making systems more adaptable to various environmental conditions. While CNNs are commonly employed as standalone feature extractors, studies investigating their enhancement through feature aggregation highlight the potential for improved generalization and robustness Su2020. This approach has been explored in both shallow machine learning models Domazetovska2021 and deep learning frameworks Liu2021, emphasizing its role in refining ESC methodologies.
@@
This study provides a systematic comparison of lightweight CNNs with stacked features versus AST under multiple data regimes. We present first a performance comparison focused primarily on CNN models utilizing stacked feature representations for ESC using the widely adopted ESC-50 dataset. We investigate the impact of different stacked feature combinations as input to CNN, including Log-Mel Spectrogram (LM), Spectral Contrast (SPC), Chroma (CH), Tonnetz (TZ), and Mel-Frequency Cepstral Coefficients (MFCCs). The evaluated combinations include LM, LM+TZ, LM+ MFCC, MFCC+TZ, LM+SPC+CH, and MFCC+GTCC +CH+LM. Additionally, we have evaluated the CNNs against the AST model, which is known for learning long-range dependencies from audio data. The findings of this study provide insights into the effectiveness of stacked feature representations in CNN-based ESC and highlight the trade-offs between CNN and AST architectures for environmental sound classification.
@@ 2 Related Work
Feature aggregation, also referred to as feature concatenation or feature-level fusion, allows models to leverage complementary information from different spectral and temporal characteristics of environmental sounds. Studies have investigated various feature aggregation approaches, including the concatenation of Log-Mel Spectrograms, Spectral Contrast, Chroma, Tonnetz, and MFCCs to enhance classification accuracy gupta2024esc; Mushtaq2020_2; sharma2019environment; zhu2018environmental. Feature selection and dimensionality reduction techniques, such as Principal Component Analysis (PCA), have also been applied to aggregated features to reduce redundancy and computational complexity Burgos2014. Deep learning has revolutionized ESC by enabling automatic feature extraction from raw audio signals or spectrogram representations. CNNs have been widely adopted for ESC due to their ability to capture local spectral-temporal features. They are commonly employed to extract features directly from spectrogram representations, which are then passed to a neural network classifier piczak2015environmental.
@@
However, to improve classification performance and develop high-efficiency AI models, some studies have explored feature aggregation as a strategy. These methods involve combining multiple spectral features, allowing models to leverage complementary information from different aspects of the audio signal. More advanced fusion strategies, such as attention-based fusion Dai2020 and multimodal feature integration, have also been proposed to enhance ESC performance sharma2019environment. However, most existing aggregation research focuses on feature aggregation as a technique within CNN models, without directly comparing its impact against standalone CNNs or ASTs. This lack of direct comparison highlights the contribution of our study, which systematically evaluates the performance of CNNs trained with single features versus CNNs with stacked features, including evaluation against the AST model. Studies have explored enhanced CNN architectures, including deep convolutional networks with regularization techniques and data augmentation to improve robustness Mushtaq2020. Nevertheless, the direct comparison between standalone CNNs and improved CNNs utilizing stacked feature fusion remains underexplored. Transformers, particularly ASTs, while not the primary focus of this study, have recently gained attention for ESC due to their ability to model long-range dependencies in audio signals. ASTs utilize self-attention mechanisms to capture contextual information across the entire spectrogram representation. Recent studies have demonstrated the effectiveness of ASTs in ESC, achieving state-of-the-art results on benchmark datasets such as ESC-50 and UrbanSound8K chen2021environmental; salamon2014dataset.
@@
However, AST models typically require extensive pre-training and large-scale datasets to achieve optimal performance, making them computationally demanding for real-time applications. Thus, larger datasets such as Audioset have been used for pretraining purposes, leveraging the AST performance in ESC applications. In this work, we compare the AST using the same datasets as the CNNs, ESC-50 and UrbanSound8K, also with a larger dataset (Audioset), against the CNN and the CNN with stacked features. Despite these advancements, there is a lack of direct comparison between CNN models with stacked feature representations, including evaluations with other architectures, particularly AST, in the context of ESC. This study aims to address this gap by conducting a comprehensive evaluation of CNNs with stacked features on the UrbanSound8K dataset, analyzing how feature aggregation influences classification performance. Additionally, we provide a comparative analysis against AST models to assess how well CNNs with aggregated features capture spectro-temporal patterns compared to self-attention-driven transformer models. This comparison serves to contextualize the efficiency of CNN-based feature stacking strategies in relation to AST architectures.
@@ 3.1 Stacked Features
To extract meaningful features from audio signals, we used the Librosa library mcfee2015librosa, which provides a wide range of methods for feature extraction in environmental sound classification. The extracted features include cepstral-based, spectral, and tonal descriptors, all of which are essential for capturing the diverse acoustic properties of environmental sounds. We investigated stacked feature inputs to leverage complementary spectral, cepstral, tonal, and auditory characteristics for enhancing CNN-based ESC performance. Inspired by the effectiveness of feature aggregation demonstrated in prior research gupta2024esc, we selected several feature combinations to provide diverse acoustic perspectives.
After extracting individual features, we aggregate them into a unified representation to serve as input to the CNN models. This process, known as feature stacking, involves combining multiple spectrogram-based features into a multi-channel image-like structure. Each extracted feature—such as MFCC, GTCC, Log-Mel Spectrogram, Chroma, Tonnetz, or Spectral Contrast—is initially represented as a two-dimensional array corresponding to the time-frequency domain. However, the shape of these arrays can differ across features. To make them compatible for stacking, we first resize all feature maps to a common dimension of 128 × 128. This resizing is achieved either through zero-padding or bilinear interpolation, depending on the original shape of the feature. Once all features share the same spatial dimensions, they are concatenated along a new axis representing channels, similar to how RGB channels are arranged in color images. For instance, when combining MFCC, GTCC, and Log-Mel Spectrogram, the resulting input tensor has a shape of 128 × 128 × 3, with each “channel” corresponding to one of the three features. In the case of MFCC + GTCC + CH + LM, the resulting stacked input has a shape of 128 × 128 × 4. This approach enables the model to learn from multiple acoustic perspectives simultaneously, capturing spectral, cepstral, and tonal characteristics of the sound. Additionally, the image-like structure of the input makes it naturally suited for convolutional neural networks, which excel at processing spatially correlated data. As shown in our experiments, this method of aggregation is both computationally efficient and effective for improving classification performance.
@@ 3.2.1 Convolutional Neural Network (CNN) architectures
CNN-1 is a lightweight architecture with four convolutional layers using 2×2 kernels and ReLU activation. The first two layers use 32 filters, and the final two use 64 filters, each followed by max pooling to reduce spatial dimensions. A global average pooling layer compresses feature maps before a 1024-unit dense layer and softmax classifier. The model is trained with the Adam optimizer (learning rate 0.001) and categorical cross-entropy loss, balancing feature learning ability and computational efficiency.
@@
CNN-2 extends the baseline model with batch normalization after all convolutional and dense layers to stabilize training and dropout (0.25) after the second and fourth convolutional layers to reduce overfitting. It retains global average pooling and the same 1024-unit dense layer with ReLU and softmax output. Both architectures were refined through empirical tuning on ESC-50, exploring variations in kernel size, filter count, dropout, and learning rate. Four convolutional layers with 2×2 kernels and 32–64 filters provided the best trade-off between accuracy and efficiency. Regularization choices in CNN-2 were inspired by prior audio classification research and validated experimentally, resulting in models that generalize well to environmental sound classification tasks. The inclusion of both CNN-1 and CNN-2 allows us to systematically examine the trade-off between architectural simplicity and regularization-driven generalization in the context of stacked feature representations.

### v2

# Evaluating CNN with Stacked Feature Representations and Audio Spectrogram Transformer Models for Sound Classification

@@ 1 Introduction
Environmental sound classification (ESC) has emerged as a crucial task in machine learning and signal processing, finding applications in diverse domains such as surveillance, smart city monitoring, fault detection in automobiles, and acoustic-based quality control in manufacturing. Unlike structured audio signals such as speech and music, environmental sounds exhibit significant variations in temporal and spectral characteristics, making classification a challenging problem. Recent advances in machine learning and deep learning have transformed the way environmental audio is analyzed  [1]. First, traditional machine learning approaches relied on handcrafted features such as Mel- Frequency Cepstral Coefficients (MFCCs), Zero-Crossing Rate (ZCR), and Spectral Contrast (SPC), among others, and their combinations [2, 3]. They were used as inputs to classifiers like Support Vector Machines (SVMs), k-Nearest Neighbors (k-NN), Decision Trees (DT), Local Binary Pattern (LPB), and Random Forest (RF)  [4, 5]. These models achieved reasonable results but often lack robustness to noise and generalization across diverse acoustic environments. With the advent of deep learning, particularly Convolutional Neural Networks (CNNs), ESC systems became capable of learning hierarchical time-frequency patterns directly from the spectrogram representations  [1]. They can efficiently extract local spectrum features, making them well-suited for multiple audio classification tasks  [6, 7, 8].
@@
However, CNNs often struggle with capturing long-range dependencies in audio sequences, which is a fundamental limitation in modeling complex environmental sounds. In contrast, transformer-based models, such as Audio Spectrogram Transformers (AST) [9], leverage self-attention mechanisms to model global dependencies in spectrogram representations. ASTs have demonstrated state-of-the-art performance in various audio classification tasks, particularly when pre-trained on large-scale datasets  [9]. However, their reliance on extensive data and computational resources poses challenges for real-world applications where labeled data is limited. Within this broader evolution, researchers have also explored feature aggregation as a complementary approach to improve learning efficiency  [10]. By combining multiple extracted features, researchers have sought to improve classification performance and reduce model complexity, making systems more adaptable to various environmental conditions. While CNNs are commonly employed as standalone feature extractors, studies investigating their enhancement through feature aggregation highlight the potential for improved generalization and robustness [11]. This approach has been explored in both shallow machine learning models [12] and deep learning frameworks [10], emphasizing its role in refining ESC methodologies.
@@
This study presents a performance comparison focused primarily on CNN models utilizing stacked feature representations for ESC using the widely adopted ESC-50 dataset. We investigate the impact of different stacked feature combinations as input to CNN, including Log-Mel Spectrogram (LM), Spectral Contrast (SPC), Chroma (CH), Tonnetz (TZ), and Mel-Frequency Cepstral Coefficients (MFCCs). The evaluated combinations include LM, LM+ TZ, LM+ MFCC, MFCC+ TZ, LM+ SPC+ CH, and MFCC+ GTCC+ CH+ LM. Additionally, we have evaluated the CNNs against the AST model, which is known for learning long-range dependencies from audio data. The findings of this study provide insights into the effectiveness of stacked feature representations in CNN-based ESC and highlight the trade-offs between CNN and AST architectures for environmental sound classification.
@@ 2 Related Work
Feature aggregation, also referred to as feature concatenation or feature-level fusion, allows models to leverage complementary information from different spectral and temporal characteristics of environmental sounds. Studies have investigated various feature aggregation approaches, including the concatenation of Log-Mel Spectrograms, Spectral Contrast, Chroma, Tonnetz, and MFCCs to enhance classification accuracy [13, 14, 15, 16]. Feature selection and dimensionality reduction techniques, such as Principal Component Analysis (PCA), have also been applied to aggregated features to reduce redundancy and computational complexity [17].
@@
Deep learning has revolutionized ESC by enabling automatic feature extraction from raw audio signals or spectrogram representations. CNNs have been widely adopted for ESC due to their ability to capture local spectral-temporal features. They are commonly employed to extract features directly from spectrogram representations, which are then passed to a neural network classifier [18]. However, to improve classification performance and develop high-efficiency AI models, some studies have explored feature aggregation as a strategy. These methods involve combining multiple spectral features, allowing models to leverage complementary information from different aspects of the audio signal. More advanced fusion strategies, such as attention-based fusion [19] and multimodal feature integration, have also been proposed to enhance ESC performance [15]. However, most existing aggregation research focuses on feature aggregation as a technique within CNN models, without directly comparing its impact against standalone CNNs or ASTs. This lack of direct comparison highlights the contribution of our study, which systematically evaluates the performance of CNNs trained with single features versus CNNs with stacked features, including evaluation against the AST model. Studies have explored enhanced CNN architectures, including deep convolutional networks with regularization techniques and data augmentation to improve robustness [20]. Nevertheless, the direct comparison between standalone CNNs and improved CNNs utilizing stacked feature fusion remains underexplored.
@@
Transformers, particularly ASTs, while not the primary focus of this study, have recently gained attention for ESC due to their ability to model long-range dependencies in audio signals. ASTs utilize self-attention mechanisms to capture contextual information across the entire spectrogram representation. Recent studies have demonstrated the effectiveness of ASTs in ESC, achieving state-of-the-art results on benchmark datasets such as ESC-50 and UrbanSound8K  [21, 22]. However, AST models typically require extensive pre-training and large-scale datasets to achieve optimal performance, making them computationally demanding for real-time applications. Thus, larger datasets such as Audioset have been used for pretraining purposes, leveraging the AST performance in ESC applications. In this work, we compare the AST using the same datasets as the CNNs, ESC-50 and UrbanSound8K, also with a larger dataset (Audioset), against the CNN and the CNN with stacked features. Despite these advancements, there is a lack of direct comparison between CNN models with stacked feature representations, including evaluations with other architectures, particularly AST, in the context of ESC. This study aims to address this gap by conducting a comprehensive evaluation of CNNs with stacked features on the UrbanSound8K dataset, analyzing how feature aggregation influences classification performance. Additionally, we provide a comparative analysis against AST models to assess how well CNNs with aggregated features capture spectro-temporal patterns compared to self-attention-driven transformer models. This comparison serves to contextualize the efficiency of CNN-based feature stacking strategies in relation to AST architectures.
@@ 3 Methodology
3.1 Features
To extract meaningful features from audio signals, we utilized the Librosa library  [23], which provides multiple methods for feature extraction in environmental sound classification. Extracted features include cepstral-based features, spectral features, and tonal features, which are essential for capturing various acoustic properties of environmental sounds. Below, we describe each extracted feature in detail.
3.1.1 Mel-Frequency Cepstral Coefficients (MFCCs)
MFCCs are derived from the log power spectrum of an audio signal  [24], using the Mel scale, which better reflects human auditory perception compared to a linear frequency scale. As illustrated in Figure1, the computation begins with the raw waveform, which is segmented into short overlapping frames. Each frame is then windowed using a Hamming window to minimize spectral leakage. The windowed frames undergo a Fast Fourier Transform (FFT) to convert the time-domain signal into the frequency domain. Subsequently, a Mel filter bank is applied to the magnitude spectrum to emphasize perceptually important frequency bands. The filtered spectral energies are converted to a logarithmic scale, and a Discrete Cosine Transform (DCT) is applied to decorrelate the features. The resulting coefficients are the MFCCs, which compactly represent the spectral envelope of the audio signal. The n-th cepstral coefficient, Cn, representing the spectral envelope features, can be represented as:
$C_{n}=\sum_{m=0}^{M-1}\log(S_{m})\,\cos\left[n\left(m-0.5\right)\frac{\pi}{M}\right],$ (1)
where $M$ is the total number of filter banks, $S_{m}$ is the spectral energy in the $m$-th frequency band (computed from the filter bank output), $n$ is the index of the cepstral coefficient (ranging from 0 to N1N-1, where $N$ is the number of coefficients extracted), $m$ is the index of the filter bank (ranging from 0 to M1M-1). $\log(S_{m})$ is the logarithm of spectral energy, used to mimic human auditory perception.
Figure 1: Block diagram of MFCC algorithm
3.1.2 Log-Mel Scale Spectrogram (LM)
A Log-Mel spectrogram [11] is a time-frequency representation of an audio signal in which the frequencies are mapped to the Mel scale and then transformed to a logarithmic scale. This representation provides a compact and perceptually relevant view of the spectral content, making it a popular choice for deep learning models. A Log-Mel spectrogram, as shown in Figure 2, is computed in a manner similar to MFCCs, but without applying the Discrete Cosine Transform (DCT). The Log-Mel Spectrogram value, Log-Mel(m), for the m-th Mel filter is:
$\mathrm{Log\text{-}Mel}(m)=\log\left(\sum_{k=0}^{N-1}\left|X(n,k)\right|^{2}H_{m}(k)+\epsilon\right),$ (2)
where $N$ is the total number of frequency bins in the spectrogram. $X(n,k)$ is the Short-Time Fourier Transform (STFT) coefficient at time frame $n$ and frequency bin $k$. $|X(n,k)|^{2}$ represents the power spectrum (magnitude squared of the STFT). $H_{m}(k)$ is the Mel filter bank response at the $m$-th filter and frequency $k$, which converts the linear frequency scale to the Mel scale. $\epsilon$ is a small constant added for numerical stability to avoid taking the logarithm of zero.
Figure 2: Block diagram of Log-Mel Scale Spectrogram algorithm
3.1.3 Chroma Features (CH)
Chroma features [11] aggregate spectral information into 12 distinct pitch classes. These features emphasize the harmonic and tonal structure of a sound. To compute them, first perform the STFT of the audio signal, then map frequency bins to 12 chroma bins corresponding to pitch classes. The Chroma feature is given by:
$C_{n}=\frac{\sum_{k\in B_{n}}\left|X(n,k)\right|^{2}}{\sum_{k=0}^{N-1}\left|X(n,k)\right|^{2}},$ (3)
where $B_{n}$ is the set of frequency bins corresponding to pitch class $n$. $|X(n,k)|^{2}$ is the STFT power spectrum.
3.1.4 Spectral Contrast (SPC)
Spectral contrast [11] measures the difference between peaks and valleys in the frequency spectrum rather than focusing on the average spectral envelope. It captures tonal richness and helps distinguish between harmonic and percussive sounds. The Spectral Contrast feature can be described as:
$SC_{m}=\frac{S_{\max,m}-S_{\min,m}}{S_{\max,m}+S_{\min,m}},$ (4)
where $S_{\max,m}$ is the maximum spectral amplitude in band $m$ and $S_{\min,m}$ is the minimum spectral amplitude in band $m$.
3.1.5 Tonnetz (TZ)
Tonnetz (Tonal Centroids) features [11], which capture the harmonic relationships within an audio signal, are computed based on chroma features. The process is illustrated as:
$T=UC,$ (5)
where $C$ is the chroma feature vector and $U$ is a transformation matrix that maps chroma into a six-dimensional tonal space.
3.1.6 Gammatone Frequency Cepstral Coefficients (GTCC)
GTCCs are similar to MFCCs but use a Gammatone filter bank instead of Mel filters [15]. GTCCs are more robust to noise and transient sounds compared to MFCCs. The computation of GTCCs is described as:
$C_{n}=\sum_{m=0}^{M-1}\log\!\left(S_{g}(m)\right)\cos\!\left[n\left(m-0.5\right)\frac{\pi}{M}\right],$ (6)
where $C_{n}$ is the $n$-th GTCC coefficient, which represents a transformed version of the Gammatone-filtered log spectrum. $M$ is the total number of Gammatone filterbanks, typically chosen based on how fine-grained the frequency analysis should be. $S_{g}(m)$ is the Gammatone power spectrum at the $m$-th frequency band, computed by applying a Gammatone filterbank to the signal’s power spectrum. $n$ is the index of the cepstral coefficient; it controls which coefficient is being computed in the Discrete Cosine Transform (DCT). $m$ is the index of the Gammatone filter; it corresponds to different frequency bands in the Gammatone filterbank. $\log(S_{g}(m))$ is the logarithm of the Gammatone power spectrum at the $m$-th band. The logarithm is applied to mimic human loudness perception, since the human ear perceives loudness logarithmically.
@@ 3.2 Stacked Features
We investigated stacked feature inputs to leverage complementary spectral, cepstral, tonal, and auditory characteristics to enhance CNN-based ESC performance. Inspired by the effectiveness of feature aggregation shown in prior research [13], we selected several combinations to provide diverse acoustic perspectives. They are presented in Table 1.
After extracting individual features, as described in Section 3.1, we aggregate them into a unified representation to serve as input to the CNN models. This process, known as feature stacking, involves combining multiple spectrogrambased features into a multi-channel image-like structure. Each extracted feature—such as MFCC, GTCC, Log-Mel Spectrogram, Chroma, Tonnetz, or Spectral Contrast—is initially represented as a two-dimensional array corresponding to the time-frequency domain. However, the shape of these arrays can differ across features. To make them compatible for stacking, we first resize all feature maps to a common dimension of 128 × 128. This resizing is achieved either through zero-padding or interpolation, depending on the original shape of the feature. Once all features share the same spatial dimensions, they are concatenated along a new axis representing channels, similar to how RGB channels are arranged in color images.
For instance, when combining MFCC, GTCC, and Log-Mel Spectrogram, the resulting input tensor has a shape of 128 × 128 × 3, with each “channel” corresponding to one of the three features. In the case of MFCC + GTCC + CH + LM, the resulting stacked input has a shape of 128 × 128 × 4. This approach enables the model to learn from multiple acoustic perspectives simultaneously, capturing spectral, cepstral, and tonal characteristics of the sound. Additionally, the image-like structure of the input makes it naturally suited for convolutional neural networks, which excel at processing spatially correlated data. As shown in our experiments, this method of aggregation is both computationally efficient and effective for improving classification performance.
Table 1: Feature Configurations and their Descriptions.
Feature Configuration Expected Contribution
LM Serves as a baseline, capturing spectral energy distribution.
LM + TZ Combines spectral and tonal features, improving discrimination between harmonic and non-harmonic sounds.
LM + SPC + CH Emphasizes distinctions between harmonic-rich, percussive, and transient sounds through combined spectral contrast and pitch class energy.
MFCC + LM Integrates cepstral and spectral information, enhancing classification across environmental sounds.
MFCC + TZ Enhances the classification of structured (tonal) and unstructured (non-tonal) sounds, such as distinguishing musical notes from random noise.
MFCC + GTCC + CH + LM Provides the most comprehensive representation, integrating cepstral, spectral, tonal, and auditory system characteristics, thus improving robustness to noise.
@@ 3.3.1 Convolutional Neural Network (CNN) architectures
Figure 3: CNN Architectures and examples of feature applied to a sample ("dog" sample from ESC-50): (a) CNN-1 architecture. (b) CNN-2 architecture. (c) Log-Mel spectrogram. (d) MFCC. (e) Chroma. (f) Spectral contrast. (g) Tonnetz. (h) GTCC.
@@
The first model (CNN-1), as we have shown in Figure3-a, is a straightforward yet effective CNN architecture consisting of four convolutional layers, followed by global average pooling and dense layers. This design is intended to efficiently extract hierarchical features while maintaining a balance between model complexity and computational efficiency. The architecture comprises four convolutional layers with 2×2 kernels, utilizing the ReLU activation function. The feature extraction starts with 32 filters in the two initial layers and progresses to 64 filters for the other two layers, enabling the network to learn increasingly complex representations. Each convolutional layer is followed by max pooling, which reduces the spatial dimensions of the feature maps, minimizing computational cost while preserving essential patterns. To further refine the extracted features, a global average pooling layer aggregates spatial features into a single feature vector, reducing dimensionality before being fed into the dense layers. The final classification is performed using a 1024-unit dense layer with ReLU activation, followed by a softmax layer for multi-class sound recognition. The model is compiled using the Adam optimizer with a learning rate of 0.001 and is trained using the categorical cross-entropy loss function to optimize performance in multi-class classification.
@@
The second model (CNN-2) as we have shown in Figure3-b, is an enhanced CNN architecture that incorporates batch normalization and dropout to improve training stability and generalization. While it follows a similar structure to CNN-1, it introduces additional regularization techniques to mitigate overfitting and optimize learning efficiency. In this architecture, Batch Normalization is applied after each convolutional and dense layer to normalize activations, improve gradient flow, and accelerate training convergence. Dropout (0.25 probability) was incorporated after the second and fourth convolutional layers to reduce overfitting by preventing co-adaptation of neurons. Global Average Pooling is utilized to reduce feature dimensionality while retaining essential spatial information before the dense layers. A 1024-unit dense layer with batch normalization and ReLU activation is followed by a softmax classifier for multi-class sound recognition.
The selection of architectural parameters for both CNN-1 and CNN-2 was guided by a combination of empirical tuning and insights from prior research. Initially, we experimented with various CNN configurations using different numbers of convolutional layers, kernel sizes, and filter counts. Through systematic testing on the ESC-50 dataset, we observed that four convolutional layers with small kernel sizes (2×2) provided a good balance between feature learning capacity and overfitting prevention. Similarly, the use of 32 and 64 filters in early and deeper layers, respectively, was found to yield strong performance while keeping the model computationally efficient. The choice of global average pooling and dense layers was inspired by common best practices in audio classification tasks and further validated through experimentation. The dropout rate (0.25) and use of batch normalization in CNN-2 were adopted from standard regularization strategies as explored in [20], [6], and similar audio classification works. We also tested different dropout rates (0.3, 0.5) and optimizer learning rates (e.g., 0.0001, 0.0005, 0.001), ultimately selecting those that led to stable convergence and the best validation performance. Thus, the final CNN configurations were the result of a combination of design inspiration from prior works and hyperparameter tuning on our development set. This process ensured that the models were both effective and generalizable to the target classification tasks.